# Tutorial on amortized optimization

### v4

@@ 2.2.2 Learning iterative semi-amortized models
[…]
The parameterization of every iterate $z_{\theta}^{i}$ can influence the final prediction $\hat{y}_{\theta}$ and thus losses on top of $\hat{y}_{\theta}$ need to consider the entire chain of computations. Differentiating through this kind of iterative procedure is referred to as backpropagation through time in sequence models and unrolled optimization [Pearlmutter and Siskind, 2008, Zhang and Lesser, 2010, Maclaurin et al., 2015b, Belanger and McCallum, 2016, Metz et al., 2017, Finn et al., 2017, Han et al., 2017, Belanger et al., 2017, Belanger, 2017, Foerster et al., 2017, Bhardwaj et al., 2020, Monga et al., 2021] when the iterates are solving an optimization problem. The term “unrolling” arises because the model computation is iterative and computing $\mathrm{D}_{\theta}[\hat{y}_{\theta}(x)]$ requires saving and differentiating the “unrolled” intermediate iterations, as in section 2.1.2. The terminology “unrolling” here emphasizes that the iterative computation produces a compute graph of operations and is likely inspired from loop unrolling in compiler optimization [Aho et al., 1986, Davidson and Jinturkar, 1995] where loop operations are inlined for efficiency and written as a single chain of repeated operations rather than an iterative computation of a single operation.
Even though $\mathrm{D}_{\theta}\hat{y}_{\theta}$ through unrolled optimization is well-defined, in practice it can be unstable because of exploding gradients [Pearlmutter, 1996, Pascanu et al., 2013, Maclaurin, 2016, Parmas et al., 2018] and inefficient for compute and memory resources because every iterate needs to be stored, as in section 2.1.2. This is why most methods using unrolled optimization for learning often only unroll through tens of iterations [Metz et al., 2017, Belanger et al., 2017, Foerster et al., 2017, Finn et al., 2017] while solving the problems from scratch may require 100k-1M+ iterations. This causes the predictions to be extremely inaccurate solutions to the optimization process and has sparked the research directions that the next section turns to that seek to make unrolled optimization more tractable.
@@
[…]
@@
In addition to truncating the iterations, other approaches attempt to improve the efficiency of learning through unrolled iterations with other approximations that retain the influence of the entire sequence of predictions on the loss [Finn et al., 2017, Nichol et al., 2018, Lorraine et al., 2020] which will be further discussed in section 2.1.2. Some optimization procedures, such as gradient descent with momentum, can also be “reversed” without needing to retain the intermediate states [Maclaurin et al., 2015b, Franceschi et al., 2017]. Real-Time Recurrent Learning (RTRL) by Williams and Zipser [1989] uses forward-mode automatic differentiation to compute unbiased gradient estimates in an online fashion. Unbiased Online Recurrent (UORO) by Tallec and Ollivier [2018] improves upon RTRL with a rank-1 approximation of the gradient of the hidden state with respect to the parameters. Silver et al. [2022] considers the directional derivative of a recurrent model along a candidate direction, which can be efficiently computed to construct a descent direction.
@@
Figure 2.4: Illustration of the penalty used in the Implicit MAML by Rajeswaran et al. [2019] in eq. 2.13. The original loss $f(y;x)$ is shown in black for a fixed context $x$ and the lighter grey colors show the impact of varying $\lambda$. This shows that the quadratic term of the penalization eventually overtakes the original loss and makes an optimum appear close to $\hat{y}_{\theta}^{0}$
@@
This section briefly summarizes Implicit MAML (iMAML) by Rajeswaran et al. [2019], which notably brings this insight to MAML. MAML methods usually only take a few gradient steps and are usually not enough to globally solve eq. 1.1, especially at the beginning of training. Rajeswaran et al. [2019] observe that adding a penalty to the objective around the initial iterate $\hat{y}^{0}_{\theta}$ makes it easy for the model to globally (!) solve the problem
@@
where the parameter $\lambda$ encourages the solution to stay close to some initial iterate. Figure 2.4 visualizes a function $f(y;x)$ in black and add penalties in grey with $\lambda\in[0,12]$ and see that a global minimum is difficult to find without adding a penalty around the initial iterate. This global solution can then be implicitly differentiated to obtain a derivative of the loss with respect to the model’s parameters without needing to unroll, as it requires significantly less computational and memory resources. Huszár [2019] further analyzes and discuses iMAML. They compare it to a Bayesian approach and observe that the insights from iMAML can transfer from gradient-based meta-learning to other amortized optimization settings.
Warning. Implicit differentiation is only useful when optimization problems are exactly solved and satisfy the conditions of the implicit function theorem in theorem 1. This is why Rajeswaran et al. [2019] needed to add a penalty to MAML’s inner optimization problem in eq. 2.13 to make the problem exactly solvable. While they showed that this works and results in significant improvements for differentiation, it comes at the expense of changing the objective to penalize the distance from the previous iterate. In other words, iMAML modifies MAML’s semi-amortized model and in general is not helpful for estimating the derivative through the original formulation of MAML. Furthermore, computing the implicit derivative by solving the linear system with the Jacobian in eq. 2.30 may be memory and compute expensive to form and estimate exactly. In practice, some methods such as Bai et al. [2019] successfully use indirect and approximation methods to solve for the system in eq. 2.30.
@@ 2.2.3 Learning with zeroth-order methods and RL
Computing the derivatives to learn $\hat{y}_{\theta}$ with a first-order method may be impossible or unstable. These problems typically arise when learning components of the model that are truly non-differentiable, or when attempting to unroll a semi-amortized model for a lot of steps. In these settings, research has successfully explored other optimizers that do not need the gradient information. These methods often consider settings that improve an objective-based loss with small local perturbations rather than differentiation. Figure 2.5 illustrates that most of these methods can be interpreted as locally perturbing the model’s prediction and updating the parameters to move towards the best perturbations.
@@
where $\sigma^{2}$ is a fixed variance. Figure 2.6 illustrates a loss function ${\mathcal{L}}$ in black and shows smoothed versions in color. They consider learning the loss with reparameterization gradients and zeroth-order evolutionary methods. Merchant et al. [2021] further build upon this for learned optimization in atomic structural optimization and study 1) clipping the values of the gradient estimator, and 2) parameter optimization with genetic algorithms.
Figure 2.6: Gaussian smoothing of a loss using eq. 2.14. The colors show different values of the variance $\sigma^{2}$ of the Gaussian. Selecting a high enough variance results in smoothing out most of the suboptimal minima.
@@
Connection to smoothing in reinforcement learning. The Gaussian smoothing of the objective ${\mathcal{L}}$ in eq. 2.14 is conceptually similar to Gaussian smoothing of the objective in reinforcement learning, i.e. the $-Q$-value, by a Gaussian policy. This happens in eq. 3.39 and is further discussed in section 3.6. The policy’s variance is typically controlled to match a target entropy Haarnoja et al. [2018] and the learning typically starts with a high variance so the policy has a broad view of the objective landscape and is then able to focus in on a optimal region of the value distribution. Amos et al. [2021] uses a fixed entropy decay schedule to explicitly control this behavior. In contrast, Metz et al. [2019a], Merchant et al. [2021] do not turn the loss into a distribution and more directly smooth the loss with a Gaussian with a fixed variance $\sigma^{2}$ that is not optimized over.
@@ 2.3 Extensions
I have intentionally scoped definition 1 to optimization problems over deterministic, unconstrained, finite-dimensional, Euclidean domains ${\mathcal{Y}}$ where the context distribution $p(x)$ remains fixed the entire training time to provide a simple mental model that allows us to focus on the core amortization principles that consistently show up between applications. This section cover extensions from this setting that may come up in practice.
@@ 2.3.1 Extensions of the domain ${\mathcal{Y}}$
A crucial extension is from optimization over deterministic vector spaces ${\mathcal{Y}}$ to stochastic optimization where ${\mathcal{Y}}$ represents a space of distributions, turning $y\in{\mathcal{Y}}$ from a vector in Euclidean space into a distribution. This comes up in section 3.6 for control, for example..
Transforming parameterized stochastic problems back into deterministic ones. This portion will mostly focus on settings that optimize over the parametric distributions. This may arise in stochastic domains for variational inference in section 3.1 and stochastic control in section 3.6. These settings optimize over a constrained parametric family of distributions parameterized by some $\lambda$, for example over a multivariate normal ${\mathcal{N}}(\mu,\Sigma)$ parameterized by $\lambda\vcentcolon=(\mu,\Sigma)$. Here, problems can be transformed back to eq. 1.1 by optimizing over the parameters with
@@
where $\lambda$ induces a distribution that the objective $f$ may use. When $\lambda$ is not an unconstrained real space, the differentiable projections discussed in section 2.3.1 could be used to transform $\lambda$ back into this form for amortization.
Figure 2.7: The Gaussian distribution can be characterized as the result of the optimization problem in eq. 2.16: constrained to the space of continuous distributions with a given mean and variance, the Gaussian distribution has the maximum entropy in comparison to every other distribution. This example parameterizes a non-Gaussian density (shown in grey) and optimizes over it using gradient steps of eq. 2.16, eventually converging to a Gaussian. An animated version is available in the repository associated with this tutorial. While the Gaussian is the known closed-form solution to this optimization problem and analytically known, more general optimization problems over densities without known solutions can also be amortized.
@@
Amortized constrained optimization problems may naturally arise, for example in the convex optimization settings in section 3.4 and for optimization over the sphere in section 4.2. Constrained optimization problems for amortization can often be represented as an extension of eq. 1.1 with
@@
where the constraints ${\mathcal{C}}$ may also depend on the context $x$. Remark 3 suggests one way of amortizing eq. 2.17 by amortizing the objective-based loss associated with the optimality conditions of the constrained problem. A budding research area studies how to more generally include constraints into the formulation. Baker [2019], Dong et al. [2020], Zamzam and Baker [2020], Pan et al. [2020], Klamkin et al. [2025], Hentenryck [2025] predict solutions to optimal power flow problems. Misra et al. [2021] learn active sets for constrained optimization. Kriváchy et al. [2020] solves constrained feasibility semi-definite programs with a fully-amortized neural network model using an objective-based loss. Donti et al. [2021] learns a fully-amortized model and optimizes an objective-based loss with additional completion and correction terms to ensure the prediction satisfies the constraints of the original problem.
${\mathcal{C}}$$x$$\pi_{\mathcal{C}}(x)$
Figure 2.8: Illustration of definition 4 showing a Euclidean projection $\pi_{\mathcal{C}}(x)$ of a point $x$ onto a set ${\mathcal{C}}$.
Differentiable projections. When the constraints are relatively simple, a differentiable projection can transform a constrained optimization problem into an unconstrained one, e.g., in reinforcement learning constrained action spaces can be transformed from the box $[-1,1]^{n}$ to the reals $\mathbb{R}^{n}$ by using the $\tanh$ to project from $\mathbb{R}^{n}$ to $[-1,1]^{n}$. Section 4.2 also uses a differentiable projection from $\mathbb{R}^{n}$ onto the sphere ${\mathcal{S}}^{n-1}$. These are illustrated in fig. 2.8 and defined as:
@@
Lack of idempotency. In linear algebra, a projection is defined to be idempotent, i.e. applying the projection twice gives the same result so that $\pi\circ\pi=\pi$. Unfortunately, projections as defined in definition 4, such as Bregman projections, are not idempotent in general and often $\pi_{\mathcal{C}}\circ\pi_{\mathcal{C}}\neq\pi_{\mathcal{C}}$ as the regularizer $\Omega$ may cause points that are already on ${\mathcal{C}}$ to move to a different position on ${\mathcal{C}}$.
Differentiable projections for constrained amortization. These can be used to cast Eq. 2.17 as the unconstrained problem eq. 1.1 by composing the objective with a projection $f\circ\pi_{\mathcal{C}}$. (Sub)differentiable projections enable gradient-based learning through the projection and is the most easily attainable when the projection has an explicit closed-form solution. For intuition, the ReLU, sigmoid, and softargmax can be interpreted as differentiable projections that solve convex optimization problems in the form of eq. 2.18. Amos [2019, §2.4.4] further discusses these and proves them using the KKT conditions:
  • The standard Euclidean projection onto the non-negative orthant +$\mathbb{R}^{n}_{+}$ is defined by
@@
  • The interior of the unit hypercube $[0,1]^{n}$ can be projected onto with the entropy-regularized optimization problem
@@
    is the binary entropy function. Eq. 2.20 has a closed-form solution given by the sigmoid or logistic function, i.e. $\pi(x)\vcentcolon=(1+e^{-x})^{-1}$.
  • The interior of the $(n-1)$-simplex defined by
[…]
@@
[…]
    where $H(y)\vcentcolon=-\sum_{i}y_{i}\log y_{i}$ is the entropy function. Eq. 2.23 has a closed-form solution given by the softargmax, i.e. $\pi(x)_{j}=e^{x_{j}}/\sum_{i}e^{x_{i}}$, which is historically referred to as the softmax.
Figure 2.9: Illustration of the second-order cone in eq. 2.24.
@@
    which is illustrated in fig. 2.9. The standard Euclidean projection is given in closed form as
@@
Manifold optimization [Absil et al., 2009, Hu et al., 2019] over non-Euclidean spaces is a thriving topic in optimization as these problems arise frequently over complex geometries in nature. One form of manifold optimization takes ${\mathcal{Y}}$ to be a Riemannian manifold rather than a real-valued spaced. This area of research has studied acceleration methods, [Duruisseaux and Leok, 2022], but less exploration has been done on amortized optimization. Section 4.2 discusses amortizing a simple constrained spherical optimization problem that can be transformed into an unconstrained Euclidean optimization problem by using projections from ambient Euclidean space. When this is not possible, a budding area of research investigates more directly including the manifold structure into the amortization process. Gao et al. [2020] amortize optimization problems over SPD spaces.

### v5

@@ 2.2.2 Learning iterative semi-amortized models
$\hat{z}^{0}_{\theta}$$\hat{z}_{\theta}^{1}$$\ldots$$\hat{z}_{\theta}^{K}$$\hat{y}_{\theta}(x)$$\mathcal{L}$}}
The parameterization of every iterate $z_{\theta}^{i}$ can influence the final prediction $\hat{y}_{\theta}$ and thus losses on top of $\hat{y}_{\theta}$ need to consider the entire chain of computations. Differentiating through this kind of iterative procedure is referred to as backpropagation through time in sequence models and unrolled optimization [Pearlmutter and Siskind, 2008, Zhang and Lesser, 2010, Maclaurin et al., 2015b, Belanger and McCallum, 2016, Metz et al., 2017, Finn et al., 2017, Han et al., 2017, Belanger et al., 2017, Belanger, 2017, Foerster et al., 2017, Bhardwaj et al., 2020, Monga et al., 2021] when the iterates are solving an optimization problem. The term “unrolling” arises because the model computation is iterative and computing $\mathrm{D}_{\theta}[\hat{y}_{\theta}(x)]$ requires saving and differentiating the “unrolled” intermediate iterations, as in section˜2.1.2. The terminology “unrolling” here emphasizes that the iterative computation produces a compute graph of operations and is likely inspired from loop unrolling in compiler optimization [Aho et al., 1986, Davidson and Jinturkar, 1995] where loop operations are inlined for efficiency and written as a single chain of repeated operations rather than an iterative computation of a single operation.
Even though $\mathrm{D}_{\theta}\hat{y}_{\theta}$ through unrolled optimization is well-defined, in practice it can be unstable because of exploding gradients [Pearlmutter, 1996, Pascanu et al., 2013, Maclaurin, 2016, Parmas et al., 2018] and inefficient for compute and memory resources because every iterate needs to be stored, as in section˜2.1.2. This is why most methods using unrolled optimization for learning often only unroll through tens of iterations [Metz et al., 2017, Belanger et al., 2017, Foerster et al., 2017, Finn et al., 2017] while solving the problems from scratch may require 100k-1M+ iterations. This causes the predictions to be extremely inaccurate solutions to the optimization process and has sparked the research directions that the next section turns to that seek to make unrolled optimization more tractable.
@@
$\hat{z}^{0}_{\theta}$$\hat{z}_{\theta}^{1}$$\ldots$$\hat{z}_{\theta}^{K-H}$$\ldots$$\hat{z}_{\theta}^{K}$$\hat{y}_{\theta}(x)$$\mathcal{L}$}}
@@
In addition to truncating the iterations, other approaches attempt to improve the efficiency of learning through unrolled iterations with other approximations that retain the influence of the entire sequence of predictions on the loss [Finn et al., 2017, Nichol et al., 2018, Lorraine et al., 2020] which will be further discussed in section˜2.1.2. Some optimization procedures, such as gradient descent with momentum, can also be “reversed” without needing to retain the intermediate states [Maclaurin et al., 2015b, Franceschi et al., 2017]. Real-Time Recurrent Learning (RTRL) by Williams and Zipser [1989] uses forward-mode automatic differentiation to compute unbiased gradient estimates in an online fashion. Unbiased Online Recurrent (UORO) by Tallec and Ollivier [2018] improves upon RTRL with a rank-1 approximation of the gradient of the hidden state with respect to the parameters. Silver et al. [2022] considers the directional derivative of a recurrent model along a candidate direction, which can be efficiently computed to construct a descent direction.
@@
Figure 2.4: Illustration of the penalty used in the Implicit MAML by Rajeswaran et al. [2019] in eq.˜2.13. The original loss $f(y;x)$ is shown in black for a fixed context $x$ and the lighter grey colors show the impact of varying $\lambda$. This shows that the quadratic term of the penalization eventually overtakes the original loss and makes an optimum appear close to $\hat{y}_{\theta}^{0}$
@@
This section briefly summarizes Implicit MAML (iMAML) by Rajeswaran et al. [2019], which notably brings this insight to MAML. MAML methods usually only take a few gradient steps and are usually not enough to globally solve eq.˜1.1, especially at the beginning of training. Rajeswaran et al. [2019] observe that adding a penalty to the objective around the initial iterate $\hat{y}^{0}_{\theta}$ makes it easy for the model to globally (!) solve the problem
@@
where the parameter $\lambda$ encourages the solution to stay close to some initial iterate. Figure˜2.4 visualizes a function $f(y;x)$ in black and add penalties in grey with $\lambda\in[0,12]$ and see that a global minimum is difficult to find without adding a penalty around the initial iterate. This global solution can then be implicitly differentiated to obtain a derivative of the loss with respect to the model’s parameters without needing to unroll, as it requires significantly less computational and memory resources. Huszár [2019] further analyzes and discuses iMAML. They compare it to a Bayesian approach and observe that the insights from iMAML can transfer from gradient-based meta-learning to other amortized optimization settings.
Warning. Implicit differentiation is only useful when optimization problems are exactly solved and satisfy the conditions of the implicit function theorem in theorem˜1. This is why Rajeswaran et al. [2019] needed to add a penalty to MAML’s inner optimization problem in eq.˜2.13 to make the problem exactly solvable. While they showed that this works and results in significant improvements for differentiation, it comes at the expense of changing the objective to penalize the distance from the previous iterate. In other words, iMAML modifies MAML’s semi-amortized model and in general is not helpful for estimating the derivative through the original formulation of MAML. Furthermore, computing the implicit derivative by solving the linear system with the Jacobian in eq.˜2.30 may be memory and compute expensive to form and estimate exactly. In practice, some methods such as Bai et al. [2019] successfully use indirect and approximation methods to solve for the system in eq.˜2.30.
@@ 2.2.3 Learning with zeroth-order methods and RL
Computing the derivatives to learn $\hat{y}_{\theta}$ with a first-order method may be impossible or unstable. These problems typically arise when learning components of the model that are truly non-differentiable, or when attempting to unroll a semi-amortized model for a lot of steps. In these settings, research has successfully explored other optimizers that do not need the gradient information. These methods often consider settings that improve an objective-based loss with small local perturbations rather than differentiation. Figure˜2.5 illustrates that most of these methods can be interpreted as locally perturbing the model’s prediction and updating the parameters to move towards the best perturbations.
@@
where $\sigma^{2}$ is a fixed variance. Figure˜2.6 illustrates a loss function $\mathcal{L}$} in black and shows smoothed versions in color. They consider learning the loss with reparameterization gradients and zeroth-order evolutionary methods. Merchant et al. [2021] further build upon this for learned optimization in atomic structural optimization and study 1) clipping the values of the gradient estimator, and 2) parameter optimization with genetic algorithms.
Figure 2.6: Gaussian smoothing of a loss using eq.˜2.14. The colors show different values of the variance $\sigma^{2}$ of the Gaussian. Selecting a high enough variance results in smoothing out most of the suboptimal minima.
@@
Connection to smoothing in reinforcement learning. The Gaussian smoothing of the objective $\mathcal{L}$} in eq.˜2.14 is conceptually similar to Gaussian smoothing of the objective in reinforcement learning, i.e. the Q-Q-value, by a Gaussian policy. This happens in eq.˜3.39 and is further discussed in section˜3.6. The policy’s variance is typically controlled to match a target entropy Haarnoja et al. [2018] and the learning typically starts with a high variance so the policy has a broad view of the objective landscape and is then able to focus in on a optimal region of the value distribution. Amos et al. [2021] uses a fixed entropy decay schedule to explicitly control this behavior. In contrast, Metz et al. [2019a], Merchant et al. [2021] do not turn the loss into a distribution and more directly smooth the loss with a Gaussian with a fixed variance $\sigma^{2}$ that is not optimized over.
@@ 2.3 Extensions
I have intentionally scoped definition˜1 to optimization problems over deterministic, unconstrained, finite-dimensional, Euclidean domains $\mathcal{Y}$} where the context distribution $p(x)$ remains fixed the entire training time to provide a simple mental model that allows us to focus on the core amortization principles that consistently show up between applications. This section cover extensions from this setting that may come up in practice.
@@ 2.3.1 Extensions of the domain $\mathcal{Y}$}
A crucial extension is from optimization over deterministic vector spaces $\mathcal{Y}$} to stochastic optimization where $\mathcal{Y}$} represents a space of distributions, turning $y\in{\mathcal{Y}}$ from a vector in Euclidean space into a distribution. This comes up in section˜3.6 for control, for example..
Transforming parameterized stochastic problems back into deterministic ones. This portion will mostly focus on settings that optimize over the parametric distributions. This may arise in stochastic domains for variational inference in section˜3.1 and stochastic control in section˜3.6. These settings optimize over a constrained parametric family of distributions parameterized by some $\lambda$, for example over a multivariate normal ${\mathcal{N}}(\mu,\Sigma)$ parameterized by $\lambda\vcentcolon=(\mu,\Sigma)$. Here, problems can be transformed back to eq.˜1.1 by optimizing over the parameters with
@@
where $\lambda$ induces a distribution that the objective $f$ may use. When $\lambda$ is not an unconstrained real space, the differentiable projections discussed in section˜2.3.1 could be used to transform $\lambda$ back into this form for amortization.
Figure 2.7: The Gaussian distribution can be characterized as the result of the optimization problem in eq.˜2.16: constrained to the space of continuous distributions with a given mean and variance, the Gaussian distribution has the maximum entropy in comparison to every other distribution. This example parameterizes a non-Gaussian density (shown in grey) and optimizes over it using gradient steps of eq.˜2.16, eventually converging to a Gaussian. An animated version is available in the repository associated with this tutorial. While the Gaussian is the known closed-form solution to this optimization problem and analytically known, more general optimization problems over densities without known solutions can also be amortized.
@@
Amortized constrained optimization problems may naturally arise, for example in the convex optimization settings in section˜3.4 and for optimization over the sphere in section˜4.2. Constrained optimization problems for amortization can often be represented as an extension of eq.˜1.1 with
@@
where the constraints $\mathcal{C}$} may also depend on the context $x$. Remark˜3 suggests one way of amortizing eq.˜2.17 by amortizing the objective-based loss associated with the optimality conditions of the constrained problem. A budding research area studies how to more generally include constraints into the formulation. Baker [2019], Dong et al. [2020], Zamzam and Baker [2020], Pan et al. [2020], Klamkin et al. [2025], Hentenryck [2025] predict solutions to optimal power flow problems. Misra et al. [2021] learn active sets for constrained optimization. Kriváchy et al. [2020] solves constrained feasibility semi-definite programs with a fully-amortized neural network model using an objective-based loss. Donti et al. [2021] learns a fully-amortized model and optimizes an objective-based loss with additional completion and correction terms to ensure the prediction satisfies the constraints of the original problem.
[…]
Figure 2.8: Illustration of definition˜4 showing a Euclidean projection $\pi_{\mathcal{C}}(x)$ of a point $x$ onto a set $\mathcal{C}$}.
Differentiable projections. When the constraints are relatively simple, a differentiable projection can transform a constrained optimization problem into an unconstrained one, e.g., in reinforcement learning constrained action spaces can be transformed from the box $1,1]^{n}$ to the reals $\mathbb{R}^{n}$ by using the $\tanh$ to project from $\mathbb{R}^{n}$ to $1,1]^{n}$. Section˜4.2 also uses a differentiable projection from $\mathbb{R}^{n}$ onto the sphere ${\mathcal{S}}^{n-1}$. These are illustrated in fig.˜2.8 and defined as:
@@
Lack of idempotency. In linear algebra, a projection is defined to be idempotent, i.e. applying the projection twice gives the same result so that $\pi\circ\pi=\pi$. Unfortunately, projections as defined in definition˜4, such as Bregman projections, are not idempotent in general and often $\pi_{\mathcal{C}}\circ\pi_{\mathcal{C}}\neq\pi_{\mathcal{C}}$ as the regularizer $\Omega$ may cause points that are already on $\mathcal{C}$} to move to a different position on $\mathcal{C}$}.
Differentiable projections for constrained amortization. These can be used to cast Eq.˜2.17 as the unconstrained problem eq.˜1.1 by composing the objective with a projection $f\circ\pi_{\mathcal{C}}$. (Sub)differentiable projections enable gradient-based learning through the projection and is the most easily attainable when the projection has an explicit closed-form solution. For intuition, the ReLU, sigmoid, and softargmax can be interpreted as differentiable projections that solve convex optimization problems in the form of eq.˜2.18. Amos [2019, §2.4.4] further discusses these and proves them using the KKT conditions:
  • The standard Euclidean projection onto the non-negative orthant $\mathbb{R}^{n}_{+}$ is defined by
@@
  • The interior of the unit hypercube $0,1]^{n}$ can be projected onto with the entropy-regularized optimization problem
@@
    is the binary entropy function. Eq.˜2.20 has a closed-form solution given by the sigmoid or logistic function, i.e. $\pi(x)\vcentcolon=(1+e^{-x})^{-1}$.
  • The interior of the (n1)(n-1)-simplex defined by
    $\Delta_{n-1}\vcentcolon=\{p\in\mathbb{R}^{n}\;|\;1^{\top}p=1\;\;{\rm and}\;\;p\geq 0\}$ (2.22)
@@
    $\pi(x)\in\operatorname*{arg\,min}_{0<y<1}\;\;-x^{\top}y-H(y)\;\;\operatorname*{s.t.}\;\;1^{\top}y=1$ (2.23)
    where $H(y)\vcentcolon=-\sum_{i}y_{i}\log y_{i}$ is the entropy function. Eq.˜2.23 has a closed-form solution given by the softargmax, i.e. $\pi(x)_{j}=e^{x_{j}}/\sum_{i}e^{x_{i}}$, which is historically referred to as the softmax.
Figure 2.9: Illustration of the second-order cone in eq.˜2.24.
@@
    which is illustrated in fig.˜2.9. The standard Euclidean projection is given in closed form as
@@
Manifold optimization [Absil et al., 2009, Hu et al., 2019] over non-Euclidean spaces is a thriving topic in optimization as these problems arise frequently over complex geometries in nature. One form of manifold optimization takes $\mathcal{Y}$} to be a Riemannian manifold rather than a real-valued spaced. This area of research has studied acceleration methods, [Duruisseaux and Leok, 2022], but less exploration has been done on amortized optimization. Section˜4.2 discusses amortizing a simple constrained spherical optimization problem that can be transformed into an unconstrained Euclidean optimization problem by using projections from ambient Euclidean space. When this is not possible, a budding area of research investigates more directly including the manifold structure into the amortization process. Gao et al. [2020] amortize optimization problems over SPD spaces.